Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-8x – 4y = -68
x + 3y = 16

Answer by unlockmath(1688)

Hello,
This
-8x – 4y = -68
x + 3y = 16
We can write as:
-8x – 4y = -68
x = 16-3y
Now substitute:
-8(16-3y) – 4y = -68
Expand out:
-128+24y-4y=-68
Add 128 and divide by 20 to get:
y=3
Plug this into the original and we have:
x=7
Make sense?
RJ
www.math-unlock.com